The star of a distant solar system explodes as a supernova. At the moment of the explosion, an resting exploration spaceship is 15 AU away from the shock wave. The shock wave of the explosion travels with 25000 km/s towards the spaceship. To save the crew, the spacecraft makes use of a special booster that uniformly accelerates at 150 m/s2 in the opposite direction. Determine if the crew manages to escape from the shock wave. (Neglect relativistic effects.)


1
Expert's answer
2021-02-12T06:12:19-0500

Calculate the time it takes for the spaceship to reach 25000 km/s:


"t=\frac{v}{a}=166667\text{ s}."

Calculate what distance the shockwave will cover during t:


"d=v_st=4.167·10^{12}\text{ m},"

or 27.85 AU.


So, only 27.85-15=12.85 AU left for the spaceship to develop the speed of 25000 km/s. Is it possible? Find the distance required for the ship to accelerate from rest to 25 000 000 m/s with acceleration of 150 m/s2:


"d=\frac{v^2}{2a}=2.08·10^{12}\text{ m},"

or 13.9 AU.

Therefore, when the shockwave will travel 27.85 AU from the star at 25 000 km/s, the speed of the spaceship will be lower than 25 000 km/s. They cannot escape.